# EEGAgent: A Unified Framework for Automated EEG Analysis Using Large Language Models

### v1

@@ Introduction
Recent advances in machine learning (ML) and deep learning (DL) have enabled more automated and scalable approaches to EEG analysis(zhou2025spiced). Modern data-driven models can learn meaningful representations directly from raw EEG signals and achieve high accuracy across tasks such as seizure detection, sleep staging, and neurological disorder diagnosis (li2025). These approaches have led to notable improvements in analysis speed, accuracy, and reproducibility. However, most existing EEG models remain task-specific and are optimized for isolated objectives. This leads to the so-called task isolation problem, wherein each model addresses only a narrow slice of EEG analysis without considering the broader context or the interdependence of concurrent tasks. In clinical scenarios, EEG analysis often involves multi-task, continuous reasoning—for example, detecting epileptiform discharges while also assessing cognitive states or sleep transitions (zhao2024systematic). Most current task-specific EEG models fail to meet this requirement, necessitating an urgent development of a unified and flexible framework to address this critical need. Fortunately, recent breakthroughs in large language models (LLMs) offer a promising pathway toward overcoming these limitations (zhang2024). With their powerful capabilities in task planning, contextual reasoning, and tool integration, LLM-powered agents can manage complex, multi-step workflows(MIR-2025-06-281). They dynamically interpret user intent and interface with external modules, making them ideal candidates for scheduling multi-task EEG analysis in a coordinated and intelligent manner.
@@
Electroencephalography (EEG) is widely used in sleep staging, emotion recognition, seizure detection, cognitive workload assessment, and other domains. These applications require methods capable of extracting informative patterns from complex, noisy brain signals. Existing approaches can be grouped into two main categories: traditional machine learning based on handcrafted features, and deep learning models that learn directly from raw data.
Traditional machine learning (ML) techniques remain popular in scenarios that demand efficiency or interpretability. Naive Bayes (NB) has been applied to seizure and drowsiness detection (Ashok2016), while Random Forests (RF) have shown robustness in tasks such as mental workload estimation and emotion recognition (messaoud2021random). Linear Discriminant Analysis (LDA) also remains effective in motor imagery-based BCI systems (Fu2020; Santos2023). Other classifiers, such as Support Vector Machines (SVM), K-Nearest Neighbors (KNN), and Logistic Regression (LR), are commonly used as baselines or as components in hybrid systems (saeidi2021neural). Deep learning (DL) models have also demonstrated strong performance in EEG analysis by automatically learning spatial and temporal features from EEG signals(zhoubrainuicl). Convolutional Neural Networks (CNNs) excel in sleep staging and motor imagery classification (efe2023cosleepnet; khademi2022transfer), while Recurrent Neural Networks (RNNs), such as LSTM and GRU, perform well in sequential tasks like seizure and emotion detection (mekruksavanich2023effective; muthunayagom2025stacked; chowdary2022emotion). More recently, Transformer-based architectures have been introduced for their ability to model long-range dependencies and support parallel computation (Sun2021; wei2023). Despite these advances, most EEG models are limited to isolated tasks and controlled environments, which reduces their applicability to real-world EEG-related tasks that often require multi-task analysis and flexible adaptation. In practice, tasks such as artifact removal, event detection, and signal classification frequently need to be handled together within a shared context. These challenges highlight the need for a unified framework that supports context-aware reasoning across diverse EEG scenarios.
@@
Recent advances in large language models (LLMs) have enabled the creation of intelligent agents that can perform complex reasoning and manage multi-step tasks. Originally designed for next-token prediction, LLMs now demonstrate impressive capabilities in language understanding, planning, and decision-making (brown2020language; radford2019LanguageMA). When augmented with modules for planning, memory, perception, and tool use, they can coordinate workflows, invoke external tools, and synthesize information to support high-level decision-making (yao2023react; shinn2023reflexion). Techniques such as Chain of Thought (CoT) (wei2022COT) and Tree of Thought (ToT) (yao2023TOT) enhance LLM reasoning via stepwise decomposition and hypothesis exploration. Planning modules help break down goals into subtasks, while memory mechanisms enable both contextual tracking and long-term knowledge retention (huang2022language; wang2023voyager; wu2022memorizingtransformers). Retrieval-Augmented Generation (RAG) further improves reasoning by allowing access to external knowledge bases such as clinical guidelines (lewis2020retrieval).

### v2

@@ Introduction
Recent advances in machine learning (ML) and deep learning (DL) have enabled more automated and scalable approaches to EEG analysis(zhou2025spiced). Modern data-driven models can learn meaningful representations directly from raw EEG signals and achieve high accuracy across tasks such as seizure detection, sleep staging, and neurological disorder diagnosis (li2025). These approaches have led to notable improvements in analysis speed, accuracy, and reproducibility. However, most existing EEG models remain task-specific and are optimized for isolated objectives. This leads to the so-called task isolation problem, wherein each model addresses only a narrow slice of EEG analysis without considering the broader context or the interdependence of concurrent tasks. In clinical scenarios, EEG analysis often involves multi-task, continuous reasoning—for example, detecting epileptiform discharges while also assessing cognitive states or sleep transitions (zhao2024systematic). Most current task-specific EEG models fail to meet this requirement, necessitating an urgent development of a unified and flexible framework to address this critical need. Fortunately, recent breakthroughs in large language models (LLMs) offer a promising pathway toward overcoming these limitations (zhang2024). With their powerful capabilities in task planning, contextual reasoning, and tool integration, LLM-powered agents can manage complex, multi-step workflows. They dynamically interpret user intent and interface with external modules, making them ideal candidates for scheduling multi-task EEG analysis in a coordinated and intelligent manner.
@@
Electroencephalography (EEG) is widely used in sleep staging, emotion recognition, seizure detection, and cognitive workload assessment, all of which require models capable of extracting informative patterns from noisy brain signals. Existing methods fall into two major categories: traditional machine learning with handcrafted features and deep learning models that learn directly from raw data.
Traditional ML approaches remain useful where efficiency or interpretability is important. Naive Bayes has been applied to seizure and drowsiness detection (Ashok2016), Random Forests to mental workload and emotion recognition (messaoud2021random), and Linear Discriminant Analysis (LDA) to motor imagery–based BCIs (Fu2020; Santos2023). Other classical classifiers such as SVM, KNN, and Logistic Regression are frequently used as baselines or in hybrid systems (saeidi2021neural). Deep learning methods have shown strong performance by automatically extracting spatial–temporal features from EEG signals (zhoubrainuicl; Wang_Zhao_Jiang_Li_Li_Pan_2024). CNNs excel in sleep staging and motor imagery (efe2023cosleepnet; khademi2022transfer), RNNs (e.g., LSTM, GRU) perform well in sequential tasks such as seizure and emotion detection (mekruksavanich2023effective; 10.1145/3746027.3755270; WANG2025130008), and Transformer-based models further improve the modeling of long-range dependencies (Sun2021; wei2023). However, most existing EEG models are designed for isolated tasks under controlled conditions, which limits their practical applicability. In real-world scenarios, tasks such as artifact removal, event detection, and signal classification often need to be performed together, underscoring the need for a unified, context-aware framework that can flexibly support multi-task EEG analysis.
Figure 1: EEGAgent framework
@@
Recent advances in large language models (LLMs) have enabled the creation of intelligent agents that can perform complex reasoning and manage multi-step tasks(wang2025cbramodcrisscrossbrainfoundation). Originally designed for next-token prediction, LLMs now demonstrate impressive capabilities in language understanding, planning, and decision-making (brown2020language; radford2019LanguageMA). When augmented with modules for planning, memory, perception, and tool use, they can coordinate workflows, invoke external tools, and synthesize information to support high-level decision-making (yao2023react; shinn2023reflexion). Techniques such as Chain of Thought (CoT) (wei2022COT) and Tree of Thought (ToT) (yao2023TOT) enhance LLM reasoning via stepwise decomposition and hypothesis exploration. Planning modules help break down goals into subtasks, while memory mechanisms enable both contextual tracking and long-term knowledge retention (huang2022language; wang2023voyager; wu2022memorizingtransformers). Retrieval-Augmented Generation (RAG) further improves reasoning by allowing access to external knowledge bases such as clinical guidelines (lewis2020retrieval).